# Diagnostic Performance of Universal-Learning Ultrasound AI Across Multiple Organs and Tasks: the UUSIC25 Challenge

### v1

@@ Abstract
IMPORTANCE While modern ultrasound systems are inherently versatile, current artificial intelligence (AI) solutions remain fragmented into single-task, organ-specific tools. This discrepancy limits workflow integration and the clinical utility of AI in general ultrasonography.
@@
DESIGN, SETTING, AND PARTICIPANTS The Universal UltraSound Image Challenge 2025 (UUSIC25), an international competition held from July to September 2025. Participants developed algorithms using a training set of 11,644 images derived from public datasets and limited private data (30%). Algorithms were evaluated on a fully independent, multi-center private test set of 2,479 images, including data from a center completely unseen during training to assess generalization.
@@
RESULTS A total of 15 valid algorithms were evaluated. The top-ranking algorithm (SMART) achieved a macro-averaged DSC of 0.854 (95% CI, 0.848-0.861) across 5 segmentation tasks and a macro-averaged AUC of 0.766 (95% CI, 0.718-0.815) for binary classification tasks. While models demonstrated high capability in anatomical segmentation (e.g., fetal head DSC: 0.942 [95% CI, 0.934-0.948]), performance variability was observed in complex diagnostic tasks subject to domain shift. Specifically, in breast cancer molecular subtyping, the top model’s performance dropped from an AUC of 0.571 (95% CI, 0.505-0.643) on the internal domain to 0.508 (95% CI, 0.469-0.544) on the unseen external center, highlighting the challenge of generalization.
@@ 1 Introduction
Ultrasound is one of the most widely used imaging modalities in medicine [1, 2]. It supports everyday practice in obstetrics and gynecology, abdominal and hepatobiliary imaging, cardiology, emergency and critical care, musculoskeletal clinics, and breast cancer screening [1, 3, 4, 5, 6]. Because scanners are relatively inexpensive, portable, and free of ionizing radiation, they are deployed as first line diagnostic tools in both tertiary hospitals and resource limited settings [7, 8, 9]. In routine workflows, the same machine and a small number of probe families are used to examine multiple organs and to answer different clinical questions within a single session, which naturally raises the question of whether a single general purpose ultrasound AI model on that scanner could safely support several organ specific examinations rather than only one narrowly defined task [10, 11, 12].
Despite rapid progress in medical AI, current ultrasound systems remain largely fragmented [13]. Most models are developed for a single organ and a single task, for example fetal biometry, thyroid nodule risk stratification, liver disease staging, or breast lesion classification [14, 15, 16, 17, 18, 19]. This “one model, one problem” paradigm means that each clinical question requires its own model, data pipeline, validation process, and maintenance plan. Such fragmentation increases deployment and life cycle costs and makes it difficult to build a coherent AI layer that can assist clinicians across departments, probes, and indications, which is precisely how ultrasound is used in real clinical environments.
In parallel, there is growing interest in general purpose or foundation models for ultrasound that learn shared representations across organs and tumors [20, 21, 22, 23]. Many malignant lesions, regardless of anatomical site, exhibit overlapping sonographic patterns, including irregular margins, heterogeneous echotexture, and abnormal posterior acoustic behavior, whereas normal organs show characteristic and reproducible structural motifs [14, 24, 25]. A model trained jointly on several organs may capture these common features and differences, potentially improving generalization and offering more interpretable decision patterns. Clarifying which groups of organs benefit most from cross organ learning, and how organ level structure and tumor level appearance jointly contribute to performance and interpretability, is essential for the rational design of future general purpose ultrasound AI systems [20, 21, 22, 26].
@@ 2.1 Challenge Design and Data
The cohort aggregates data from 9 public repositories (n=10,010) and internal retrospective collections from 3 partner hospitals (n=6,011), covering seven anatomical regions: breast, thyroid, liver, kidney, fetal head, cardiac, and appendix. The private in-house data were sampled from Zhejiang Cancer Hospital and Hangzhou First People’s Hospital in China, and the Netherlands Cancer Institute (NKI). This study was approved by the Medical Ethics Committee of Zhejiang Cancer Hospital (Approval No. IRB-2024-494 (IIT)). The requirement for informed consent was waived due to the retrospective nature of the study, and all images were de-identified and converted to a standardized format.
@@ 2.2 Data Partitioning Strategy
Figure 1: Study Flow Diagram. Data collection, source attribution, and stratification logic. The diagram illustrates the explicit separation of data streams: public datasets were utilized exclusively for training to promote generalization (n=10,010), while internal private data were stratified across all sets (n=5,499). Data from the external center (NKI, n=512) served as a strictly held-out test set. Data Sources: Public datasets included BUSI [27], BUSIS [28], and BUS-BRA [29] for breast; DDTI [30] for thyroid; Fatty-Liver [31] for liver; KidneyUS [32] for kidney; Fetal HC [33] for fetal head; CAMUS [34] for cardiac; and the Appendix Dataset [35]. Abbreviations: Internal centers in China include Zhejiang Cancer Hospital and Hangzhou First People’s Hospital. The external center in the Netherlands refers to the Netherlands Cancer Institute (NKI).
@@ 3 Results
3.1 Multidimensional Performance Analysis
@@
Versatility Across Organs (Figure 2A and Figure 2B). The evaluation revealed distinct performance profiles across anatomical regions. Top-performing algorithms maintained high segmentation capability (DSC >0.85>0.85) across Breast, Cardiac, Fetal Head, and Kidney datasets. However, performance consistently dipped for the Thyroid task, reflecting the intrinsic difficulty of delineating isoechoic nodules compared to structures with distinct boundaries. In classification tasks, while general-purpose models demonstrated broad applicability, trade-offs were observed; algorithms optimized for breast malignancy screening often showed reduced sensitivity in non-cancerous tasks such as appendicitis diagnosis.
@@
Efficiency-Accuracy Trade-off (Figure 2C). Clinical translation requires models that are not only accurate but also deployable. We identified a distinct trade-off between diagnostic accuracy and computational resource consumption. The winning algorithm (SMART) achieved an optimal balance, delivering top-tier diagnostic accuracy with a moderate inference time (36.03s) and a low memory footprint (0.59GB), making it suitable for standard clinical consoles. In contrast, while the fastest algorithm achieved inference speeds of 7.14s, it incurred a massive memory cost (12.41GB), limiting its utility on portable ultrasound devices. Most Transformer-based solutions converged in the center of the efficiency spectrum, suggesting an emerging architectural standard for medical foundation models.
@@
Discriminative Ability (Figure 2D-E). ROC analysis highlighted the robustness of general-purpose models in handling disparate visual tasks. In the focal task of breast malignancy classification, the top models achieved high specificity in the low false-positive rate region, a critical trait for minimizing unnecessary biopsies. Similarly, in texture-based fatty liver diagnosis, these architectures effectively captured diffuse echotexture changes with AUCs exceeding 0.85. However, performance was notably lower for complex diagnostic tasks subject to high inter-observer variability, such as appendicitis and molecular subtyping (see eFigure 3 in Supplement), underscoring current limits in extracting sub-visual biomarkers. Furthermore, visual inspection of segmentation probability maps confirmed that top-performing models maintained high boundary adherence across varying morphological complexities, from large fetal heads to irregular breast tumors (eFigure 2 in Supplement).
Figure 2: Comprehensive Evaluation of Diagnostic Performance and Efficiency. (A) Segmentation performance profiles (Dice Similarity Coefficient, DSC) across five anatomical regions for the top-15 algorithms. (B) Classification performance profiles (Area Under the Curve, AUC) across diverse diagnostic tasks. Note that while top models show consistent segmentation capability in (A), classification performance in (B) varies significantly by task difficulty. (C) Efficiency-Accuracy trade-off. The y-axis represents the composite performance score (higher is better), and the x-axis shows inference time (left is better). Marker size indicates GPU memory usage. A distinct trade-off is visible, with top-ranking algorithms occupying the optimal zone (high accuracy, low resource usage). (D-E) Receiver Operating Characteristic (ROC) curves for two clinically distinct tasks: Breast Malignancy (focal lesion) and Fatty Liver (diffuse texture). Top models maintain high discrimination (AUC > 0.83) in both scenarios. Additional visualizations are available in eFigure 2 and eFigure 3 in the Supplement.
@@ 3.2 Challenge Participation and Methods Overview
The initiative attracted 67 registered teams from Asia, Europe, and North America, integrating data sources spanning multiple continents to test cross-population generalization (eFigure 1 in Supplement).
A total of 15 valid algorithms were evaluated. In terms of methodology, there was a decisive shift toward Transformer-based architectures designed to capture long-range contextual information, distinct from traditional Convolutional Neural Networks (CNNs). Regarding integration strategies, top performers favored unified or semi-shared pipelines over decoupled models, solving segmentation and classification within a single representational structure. Detailed technical descriptions and architecture diagrams for all valid algorithms are provided in the eMethods in the Supplement.
@@ 3.4 Robustness to Domain Shifts: The Generalization Gap
Performance Collapse on Unseen Data. While algorithms achieved moderate success on internal data, we observed a consistent performance degradation on the external dataset. As detailed in Table 2, the classification-specialized algorithm (Rank 3) achieved the highest internal performance with an AUC of 0.612 (95% CI, 0.545-0.674), significantly outperforming random guessing. However, this model exhibited the most severe generalization gap (-0.138), dropping to an AUC of 0.474 (95% CI, 0.439-0.506) on the external cohort—statistically indistinguishable from chance.
@@ 4.1 Principal Findings and Clinical Implications
Third, deployment efficiency is attainable. We identified algorithms operating within the "Pareto optimal" zone, delivering high accuracy with low latency (<40<40s) and memory footprints suitable for standard consoles. However, the path to adoption faces regulatory hurdles; current frameworks (e.g., FDA) designed for single-indication devices must evolve to validate multi-task systems, likely requiring novel change control plans to ensure safety across all supported organs.
@@ 4.2 Limitations
Several limitations must be acknowledged. First, the study was retrospective and limited to B-mode ultrasound across seven organs, excluding modalities like Color Doppler or elastography. Second, the tasks were simplified to segmentation and classification, omitting complex clinical workflows such as biometric measurement, report generation, and patient history integration. Third, we did not perform a direct human-versus-model comparison or assess downstream clinical impact on patient outcomes. Finally, efficiency metrics were measured in a server environment, which may not fully reflect performance on resource-constrained edge devices. Future work should prioritize prospective validation and the expansion of tasks to dynamic video analysis. To guide the translation of these findings into practice, we provide a comprehensive roadmap for architecture design, data strategy, and evaluation standards in the Supplemental Discussion (including eTable 5) in Supplement
@@ Acknowledgments
This work was supported by the [Name of Foundation/Agency] under Grant No. XXX. We thank the organizing committee of the 28th International Conference on Medical Image Computing and Computer Assisted Intervention (MICCAI 2025) for hosting the challenge held as part of the study reported in this article, the collaborators, and the funding agencies.

### v2

@@ Abstract
IMPORTANCE Modern ultrasound systems are universal, economically cost-effective, and portable diagnostic tools capable of imaging the entire body. However, current artificial intelligence (AI) solutions remain fragmented into single-task, organ-specific tools. This critical gap between hardware versatility and software specificity limits workflow integration and the clinical utility of AI in general ultrasonography.
@@
DESIGN, SETTING, AND PARTICIPANTS The Universal UltraSound Image Challenge 2025 (UUSIC25), an international competition was held from July to September 2025. Participants developed algorithms using a training set of 11,644 images aggregated from 12 discrete sources (9 public datasets and 3 partner hospitals). Algorithms were evaluated on a fully independent, multi-center private test set of 2,479 images, composed of held-out internal samples and a cohort from an external center completely unseen during training to assess generalization.
@@
RESULTS A total of 15 valid algorithms were evaluated. The top-ranking algorithm (SMART) achieved a macro-averaged DSC of 0.854 (95% CI, 0.848-0.861) across 5 segmentation tasks and a macro-averaged AUC of 0.766 (95% CI, 0.718-0.815) for binary classification tasks. While models demonstrated high capability in anatomical segmentation (e.g., fetal head DSC: 0.942 [95% CI, 0.934-0.948]), performance variability was observed in complex diagnostic tasks subject to domain shift. Specifically, in breast cancer molecular subtyping, the top model yielded a limited baseline AUC of 0.571 (95% CI, 0.505-0.643) on the internal domain, which further dropped to 0.508 (95% CI, 0.469-0.544) on the unseen external center, highlighting both the inherent difficulty of this diagnostic task and the challenge of generalization.
@@ 1 Introduction
In clinical practice, a single modern ultrasound system, equipped with a versatile family of probes, serves as a "universal" diagnostic tool capable of imaging the entire human body—from the thyroid and breast to the liver and fetus [van2024flexible, kaselitz2025point, arnaout2021ensemble, houssami2025breast, conner2025perspectives, baloescu2025artificial]. However, the current landscape of artificial intelligence (AI) in ultrasonography stands in stark contrast to this hardware versatility. Existing AI solutions are predominantly fragmented, developed as "single-organ, single-task" tools [cid2021effect, maganti2025cardiopulmonary, nafade2024value, gao2024explainable, zhang2023predicting]. This discrepancy creates a fundamental utility gap: while the sonographer seamlessly switches probes to scan different organs within a single session [wang2020anatomy, frija2021improve, vega2025overcoming], the supporting AI software cannot.
This "one model, one problem" paradigm imposes severe limitations on clinical integration [fiorentino2023review]. It requires hospitals to deploy, validate, and maintain dozens of isolated algorithms, leading to prohibitive computational costs and complex user interactions [peng2021deep, qian2021prospective, slimani2023fetal, tong2023integration, survarachakan2022deep, yan2024domain]. To bridge this gap, we posed a fundamental question: Can a single, general-purpose deep learning model replicate the versatility of a sonographer? To answer this, we organized the Universal UltraSound Image Challenge 2025 (UUSIC25) inviting the global community to develop unified architectures capable of performing segmentation and classification across seven anatomical regions simultaneously.
In parallel, there is growing interest in general purpose or foundation models for ultrasound that learn shared representations across organs and tumors [jiao2024usfm, jiang2025pretraining, lin2025orchestration, chen2024multi]. Many malignant lesions, regardless of anatomical site, exhibit overlapping sonographic patterns, including irregular margins, heterogeneous echotexture, and abnormal posterior acoustic behavior, whereas normal organs show characteristic and reproducible structural motifs [peng2021deep, chen2023thyroid, pawlowska2024curated]. A model trained jointly on several organs may capture these common features and differences, potentially improving generalization and offering more interpretable decision patterns. Clarifying which groups of organs benefit most from cross organ learning, and how organ level structure and tumor level appearance jointly contribute to performance and interpretability, is essential for the rational design of future general purpose ultrasound AI systems [jiao2024usfm, jiang2025pretraining, lin2025orchestration, cerrolaza2019computational].
@@ 2.1 Challenge Design and Data
The cohort aggregates data from 9 public repositories (n=10,010) and internal retrospective collections from 3 partner hospitals (n=6,011), covering seven anatomical regions: breast, thyroid, liver, kidney, fetal head, cardiac, and appendix. The private in-house data were sampled from Zhejiang Cancer Hospital and Hangzhou First People’s Hospital in China, and the Netherlands Cancer Institute (NKI). This study was approved by the Medical Ethics Committee of Zhejiang Cancer Hospital (Approval No. IRB-2024-494 (IIT)), Netherlands Cancer Institute (Approval No. IRBd21-058). The requirement for informed consent was waived due to the retrospective nature of the study, and all images were de-identified and converted to a standardized format.
@@ 2.2 Data Partitioning Strategy
Figure 1: Study Flow Diagram. Data collection, source attribution, and stratification logic. The diagram illustrates the explicit separation of data streams: public datasets were utilized exclusively for training to promote generalization (n=10,010), while internal private data were stratified across all sets (n=5,499). Data from the external center (NKI, n=512) served as a strictly held-out test set. Data Sources: Public datasets included BUSI [al2020dataset], BUSIS [zhang2022busis], and BUS-BRA [gomez2024bus] for breast; DDTI [pedraza2015open] for thyroid; Fatty-Liver [byra2018transfer] for liver; KidneyUS [singla2023open] for kidney; Fetal HC [van2018automated] for fetal head; CAMUS [leclerc2019deep] for cardiac; and the Appendix Dataset [marcinkevics_2023_7711412]. Abbreviations: Internal centers in China include Zhejiang Cancer Hospital and Hangzhou First People’s Hospital. The external center in the Netherlands refers to the Netherlands Cancer Institute (NKI).
@@ 3 Results
3.1 Overview of Generalist Capabilities
@@
Clinical Versatility: Can One Model Handle Diverse Anatomies? (Figure 2B. The evaluation revealed distinct performance profiles across anatomical regions. Top-performing algorithms maintained high segmentation capability (DSC >0.85>0.85) across Breast, Cardiac, Fetal Head, and Kidney datasets. However, performance consistently dipped for the Thyroid task, reflecting the intrinsic difficulty of delineating isoechoic nodules compared to structures with distinct boundaries. In classification tasks, while general-purpose models demonstrated broad applicability, trade-offs were observed; algorithms optimized for breast malignancy screening often showed reduced sensitivity in non-cancerous tasks such as appendicitis diagnosis.
@@
Deployment Feasibility: High Performance with Minimal Resource Cost (Figure 2B). Clinical translation requires models that are not only accurate but also deployable. We identified a distinct trade-off between diagnostic accuracy and computational resource consumption. The winning algorithm (SMART) achieved an optimal balance, delivering top-tier diagnostic accuracy with a moderate inference time (36.03s) and a low memory footprint (0.59GB), making it suitable for standard clinical consoles. In contrast, while the fastest algorithm achieved inference speeds of 7.14s, it incurred a massive memory cost (12.41GB), limiting its utility on portable ultrasound devices. Most Transformer-based solutions converged in the center of the efficiency spectrum, suggesting an emerging architectural standard for medical foundation models.
@@
Diagnostic Precision: Does Generalization Compromise Accuracy? (Figure 2D). ROC analysis highlighted the robustness of general-purpose models in handling disparate visual tasks. In the focal task of breast malignancy classification, the top models achieved high specificity in the low false-positive rate region, a critical trait for minimizing unnecessary biopsies. Similarly, in texture-based fatty liver diagnosis, these architectures effectively captured diffuse echotexture changes with AUCs exceeding 0.85. However, performance was notably lower for complex diagnostic tasks subject to high inter-observer variability, such as appendicitis and molecular subtyping (see eFigure 2 in Supplement), underscoring current limits in extracting sub-visual biomarkers. Furthermore, visual inspection of segmentation probability maps confirmed that top-performing models maintained high boundary adherence across varying morphological complexities, from large fetal heads to irregular breast tumors (eFigure 1 in Supplement).
Figure 2: Global Landscape, Methodological Diversity, and Performance Benchmarking of the UUSIC25 Challenge. (a) Global Participation & Data Integration Network. The map visualizes the geographical distribution of participating research teams and the origins of clinical datasets, spanning five continents. This broad geographical coverage was designed to mitigate site-specific biases often found in single-center datasets and highlights the diversity of the patient population used for model training and evaluation. (b) Methodological Configuration Matrix. A summary of architectural choices for the top-10 teams, categorizing backbones (Transformer vs. CNN), prior integration (implicit queries vs. explicit tags), and training pipelines. Note: This panel provides a high-level overview; detailed technical descriptions and architecture diagrams for all valid algorithms are provided in the eMethods in the Supplement. (c) Multi-Organ Diagnostic Versatility and Efficiency Pane. Radar charts displaying segmentation (Dice Similarity Coefficient, DSC) and classification (Area Under Curve, AUC) performance across anatomical regions. The winning model (SMART) demonstrates consistent coverage across tasks. The scatter plot from the Efficiency Pane (bubble size = GPU memory) positions algorithms based on inference speed and aggregate performance score, identifying the "Pareto optimal" zone for clinical deployment. (d) Diagnostic Precision (ROC Curves). Receiver Operating Characteristic curves for the two most distinct clinical tasks: Breast Malignancy (focal lesion) and Fatty Liver (diffuse texture), evaluated on the private test set.
@@ 3.2 Challenge Participation and Methods Overview
The initiative attracted 74 registered teams from Asia, Europe, Africa and America, integrating data sources spanning multiple continents to test cross-population generalization (Figure 2a). A total of 15 valid algorithms were evaluated.
As summarized in Figure 2b, there was a decisive shift toward Transformer-based architectures designed to capture long-range contextual information, distinct from traditional Convolutional Neural Networks (CNNs). Regarding integration strategies, top performers favored unified or semi-shared pipelines over decoupled models. Detailed technical descriptions, specific hyper-parameter settings, and architecture diagrams for each team are comprehensively detailed in the eMethods in the Supplement.
@@ 3.4 Robustness to Domain Shifts: The Generalization Gap
Performance Collapse on Unseen Data. While algorithms achieved moderate success on internal data, we observed a consistent performance degradation on the external dataset. As detailed in Table 2, the classification-specialized algorithm (Rank 3) achieved the highest internal performance with an AUC of 0.612 (95% CI, 0.545-0.674), significantly outperforming random guessing, though the absolute performance remains modest, reflecting the challenge of extracting molecular features from B-mode ultrasound. However, this model exhibited the most severe generalization gap (-0.138), dropping to an AUC of 0.474 (95% CI, 0.439-0.506) on the external cohort—statistically indistinguishable from chance.
@@
3.5 Universal vs. Specialist: Does Versatility Compromise Accuracy?
A critical concern in developing general-purpose AI is the potential for "negative transfer," where learning multiple tasks degrades performance on individual tasks. We compared our top-ranking universal model (SMART) against state-of-the-art (SOTA) specialist models dedicated to single organs.
Notably, in the breast lesion segmentation task, SMART achieved a Dice Similarity Coefficient (DSC) of 0.854 (95% CI, 0.848-0.861). This performance is highly competitive with specialized architectures, such as the Boundary-rendering network [huang2022boundary], which reported a DSC of 0.854 on similar cohorts. This suggests that universal architectures can effectively learn shared sonographic representations—such as tumor boundaries—without compromising organ-specific accuracy.
Beyond breast imaging, SMART demonstrated robust generalization in fetal and abdominal tasks. For fetal head segmentation, it achieved a DSC of 0.942 (95% CI, 0.934-0.948). While slightly below the 0.96–0.98 range reported in idealized single-task benchmarks [torres2022review], this result is remarkable given the model’s multi-task constraint and the challenging, heterogeneous nature of the UUSIC25 test set. Similarly, in kidney segmentation, the model attained a DSC of 0.906 (95% CI, 0.892-0.919), approaching the inter-observer variability of human experts (Dice $\approx$ 0.935) as reported in recent comparative studies [valente2023comparative]. Collectively, these results indicate that a single "Generalist AI" can rival specialist models across diverse anatomies, effectively debunking the myth that versatility must come at the cost of precision.
@@ 4.1 Principal Findings and Clinical Implications
Third, deployment efficiency is attainable. We identified algorithms operating within the "Pareto optimal" zone, delivering high accuracy with low latency (<40<40s) and memory footprints suitable for standard consoles. To demonstrate this translational potential, we developed a prototype software interface (eShowcase in Supplement) incorporating the top-performing model. This prototype illustrates how the single-network architecture can automatically adapt to different organ inputs without manual configuration, providing real-time segmentation and classification results to assist sonographers. However, the path to adoption faces regulatory hurdles; current frameworks (e.g., FDA) designed for single-indication devices must evolve to validate multi-task systems, likely requiring novel change control plans to ensure safety across all supported organs.
@@ 4.2 Limitations
Several limitations must be acknowledged. First, the study was retrospective and limited to B-mode ultrasound across seven organs, excluding modalities like Color Doppler or elastography. Second, the tasks were simplified to segmentation and classification, omitting complex clinical workflows such as biometric measurement, report generation, patient history integration, and longitudinal risk assessment [wang2025mammo, wang2025predicting]. Third, we did not perform a direct human-versus-model comparison or assess downstream clinical impact on patient outcomes. Finally, efficiency metrics were measured in a server environment, which may not fully reflect performance on resource-constrained edge devices. Future work should prioritize prospective validation and the expansion of tasks to dynamic video analysis. To guide the translation of these findings into practice, we provide a comprehensive roadmap for architecture design, data strategy, and evaluation standards in the Supplemental Discussion (including eTable 5) in Supplement
4.3 Clinical Implications and Democratization of Healthcare
The transition from multiple specialist models to a single generalist model offers a quantifiable economic advantage. In a standard clinical workflow covering these 7 anatomical regions, deploying 7 separate SOTA models would typically require loading distinct weights, potentially consuming over 10 GB of VRAM. In contrast, the winning SMART algorithm requires only 0.59 GB of GPU memory. This represents a reduction of approximately 94% in computational overhead.
This efficiency has profound implications for global health equity. By lowering the hardware prerequisites, advanced AI diagnostics are no longer confined to high-end workstations in tertiary hospitals. They can be deployed on portable, hand-held, or older generation ultrasound devices common in low-and-middle-income countries (LMICs). This capability provides a tangible opportunity to democratize access to expert-level diagnostics, allowing limited medical resources to serve broader populations in resource-constrained settings.
@@ Acknowledgments
This work was supported by the Science and Technology Development Fund, Macau SAR (File no. 0004/2025/ASJ) under the FDCT-FAPESP Joint Funding Scheme; the Shenzhen Medical Research Fund (Grant No. D2501013); and the Macao Polytechnic University Grant (Grant No. RP/FCA-17/2025). We thank the organizing committee of the 28th International Conference on Medical Image Computing and Computer Assisted Intervention (MICCAI 2025) for hosting the challenge held as part of the study reported in this article, as well as our collaborators.